xx Norms

Norms

Frank Herrmann <herrmann@aei.mpg.de> Erik Schnetter <Erik.Schnetter@aei.mpg.de>

March 17 2005

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11gridfunctionsx1st 10Scope: privatex01  STRING

11Description: Set of grid functions to compute norms for 2nd order variables

11 Range 10 x01 Default: (none)
11.*  10List of grid function names

11

11gridfunctionsx2nd 10Scope: privatex01  STRING

11Description: Set of grid functions to compute norms for 2nd order variables

11 Range 10 x01 Default: (none)
11.*  10List of grid function names

11

11maxxnrxvars 10Scope: privatex01  INT

11Description: maximum number of variables for which norms to compute

11 Range 10 x01 Default: 100
110:*  10positive

11

11verbose  10Scope: privatex01  INT

11Description: verbosity of output

11 Range 10 x01 Default: (none)
110:9  10higher number means more verbose output

11

7 Interfaces

General Implements: norms

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 handlesxgroup 10  x01  compact110
11                 10  sumxhandle  x01 description11handles for reduction operators
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11 varinfo 10 x01 compact110
11                 10 varindicesx1st x01 description11variable info - which vars to compute norms for
11                 10 varindicesx2nd x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY
11 10 x01 size11MAXxNRxVARS
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11 nrdecvarxgr 10  x01  compact110
11                 10  nr1stvars  x01 description11number of vars to compute for 1st & 2nd order vars
11                 10  nr2ndvars  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11 bgvalsxgroup 10  x01  compact110
11                 10  bgvalsx1st  x01 description11background value for variable - these get subtracted
11                 10  bgvalsx2nd  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY
11 10 x01 size11MAXxNRxVARS
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 normxgroup 10  x01  compact110
11                 10  norm  x01 description11computed norm
11                 10  normxu  x01 dimensions110
11                 10  normxv  x01 distribution11CONSTANT
11                 10  normxgrad  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 diffxtermxgroup 10  x01  compact110
11                 10  diffxterm  x01 description11diff terms for norm
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn CactusNumerical/Norms. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage

11Always:  11x
11varinfo nrdecvarxgr bgvalsxgroup11x

11x  11x

11

Scheduled Functions

CCTKxWRAGH x normsxsetupxvars x check normxtype and setup the vars info x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxANALYSIS x normsxcomputexnorms x compute the norm x

11x11Language:11c
11 x 11Storage: 11diffxtermxgroup
11 x 11x  11normxgroup
11 x 11Triggers: 11normxgroup
11 x 11Type: 11function